# Scaling limits of critical FK-decorated random planar maps with $q=4$

### v1

@@ 3.1 Exact expression of the spectral density
We first recall a few properties of the resolvent function $W$ introduced in (2.22), which are part of the so-called one-cut lemma proved in [borot2012more, Section 6] using combinatorial arguments. The basic main object we use to derive the solution of (2.23) is the spectral density $\rho$ introduced in [borot2012loop], defined as follows:
@@
It is known that $\rho$ is a positive and continuous function defined on the cut $[a,b]$ that vanishes at the endpoints $a$ and $b$, and it can be deduced from [borot2012more, Equation 6.12] that $\rho$ is differentiable on (a,b)(a,b). [^7] In particular, these properties will justify the applications of the Sokhotsi–Plemelj formula below. Alternatively, we may simply assume that the Sokhotsi–Plemelj formula can be applied to $\rho$, and then after deriving our expression for $\rho$, justify this fact a posteriori by uniqueness. Recall that (2.22) implies $W(z)\sim\frac{1}{z}$ as $z\to\infty$. The Cauchy integral formula then implies that
@@
By the Sokhotski-Plemelj formula [MandalB.N2011Asie, Theorem 3.2.3], we have
@@
Then $\Phi$ is a holomorphic function in ${\mathbb{C}^{*}}\setminus[-1,1]$ whose spectral density along the cut [1,1][-1,1] is $\hat{\rho}$. By the Sokhotski-Plemelj formula [MandalB.N2011Asie, Theorem 3.2.3],
@@ 3.2 Expression and asymptotics of $F_{\ell}$
The first two claims of the above proposition are nothing but the statement of 1.4. Noting from 3.3 that the right endpoint of the cut is $b=2\sqrt{2}$, the asymptotics (3.21) match the form given by e.g. [borot2012loop, Equation (5.1)], with the additional logarithmic correction that is expected from the critical case ($n=2$).
@@ 4.1 Law of a typical cluster
By Sheffield’s bijection, we may encode $\mathfrak{t}$ as a word $w$ with $\overline{w}=\varnothing$. More precisely, Sheffield’s bijection is between the word $w$ and the map $\mathfrak{t}^{\prime}$ (in one-to-one correspondence with $\mathfrak{t}$) which is obtained from $\mathfrak{t}$ by joining all the vertices on the external face of $\mathfrak{t}$ to a single extra vertex. See Section 2.4. On the other hand, conditioned on $X(0)=\mathsf{F}$ and $X(\varphi(0))=\mathsf{h}$, the envelope $\mathfrak{e}(0)$ at 0 is encoded by the $\mathsf{F}$-excursion $E:=X(\varphi(0))\cdots X(0)$. By Lemma 2.3, the probability we are after is
@@ 4.3 Exponents for the typical loop and cluster
Our next result deals with the length $|\mathfrak{L}(0)|$ of the typical loop $\mathfrak{L}(0)$, defined as the number of triangles it crosses. This is the analogue, for the critical FK model, of the first part of [berestycki2017critical, Theorem 1.1]. Note that the latter also contains a statement about the area inside $\mathfrak{L}(0)$. We should also be able to obtain such a statement from an asymptotic that we will derive in the next section, namely (5.22). However this would require some more work, and since we will not need it in the proof of our main result, we will not pursue this direction.
@@ 5.1 The step distribution of the reduced walk
We follow the stategy of [doney1982exact, Proof of Theorem 1]. First, we relate the generating function $G(z)=\mathbb{E}[z^{\xi}]$ of $\xi$ and $r$ of (5.1). Decomposing according to the first jump, we have for all $s\in(0,1)$,
@@
One could deduce (5.9) from the fact that $\chi=2$ in Sheffield’s notation, see [SheffieldScott2016QGAI, Section 3.1]. Note, however, that this would require the difficult implication of [SheffieldScott2016QGAI, Lemma 3.1].
@@
We remark that 5.2 entails $F_{\xi}(\lambda)=1+o(\lambda)$ as $\lambda\to 0$. Since $\xi\geq-1$, we may use [bingham1974asymptotic, Section 2] to deduce that $\xi$ has finite first moment and that $\mathbb{E}[\xi]=0$, which proves (5.9).
We now turn to (5.10). Using [bingham1989regular, Corollary 8.1.7] and the Laplace transform asymptotics in 5.2, we get for free that
@@
with $a_{1}=2/\pi^{2}$. The asymptotics of the distribution function in (5.10) are more delicate and require some de Haan theory, because our case is critical (which amounts to $\alpha=1$ in the notation of [bingham1989regular, Corollary 8.1.7]). In this case one needs to check an extra condition, which is recalled in the discussion following [bingham1989regular, Corollary 8.1.7]. Namely, one needs to check that, if we set $\Phi(y):=y(1-{F}_{\xi}(1/y))$, then for all $r\geq 1$,
@@ 5.2 Scaling limit of the reduced walks
Recall that $\xi$ is the step distribution of $\mathtt{h}$ and $\mathtt{c}$. The asymptotics 5.2 for the Laplace transform of $\xi$ allow us to obtain the following scaling limit result for $\mathtt{h}$ (by symmetry, the same statement holds for $\mathtt{c}$). Recall the constant $a_{1}$ from Lemma 5.1.
@@
This Laplace transform is that of a $1$-stable random variable, see e.g. [sato1999levy, Chapter 3, Theorem 14.11]. Tracing the constants (see Lemma 5.1), this concludes the proof of 5.5. ∎
@@
We now argue that 5.6 can be seen as a scaling limit result for the primal and dual boundary length processes, when exploring towards a marked point in the infinite FK(4) random planar map. The lazy reduced walks $h^{\leftarrow}$ and $c^{\leftarrow}$ are indeed counting respectively the primal and dual boundary lengths of the component containing the triangle at 0 at each time in Sheffield’s exploration of the map, seen from infinity (see Figure 8). Whenever the lazy walk encounters a symbol $\mathsf{F}$, it jumps by an amount equal to the perimeter of the corresponding envelope, which gets disconnected from the triangle at 0. By [AHPS, lqgcle4], the limiting pair $(\zeta,\zeta^{\prime})$ in 5.6 has the same law as the left/right quantum boundary lengths at time $1$ in a uniform $\mathrm{CLE}_{4}$-exploration of a critical $\gamma=2$ quantum cone towards a Liouville-typical point. [^8] Technically the results of [AHPS] are stated for quantum discs, but they are expected to extend to other types of quantum surfaces. Therefore we may interpret 5.6 as a scaling limit result for the left(=primal) and right(=dual) boundary length process in the critical FK map, to their (conjectural) continuum analogues. A convenient way to see this is to embed the portion of the map from -$\infty$ to 0 in the half-plane, as in Figure 8.
@@ 5.3 Joint law with the duration
Moreover, for $\lambda\in[\lambda_{n+1},\lambda_{n})$ we have $\psi(\lambda)\in[\psi(\lambda_{n+1}),\psi(\lambda_{n})]$ because $\eta$ is positive, so that the asymptotic (5.23) holds with $\lambda$ in place of $\lambda_{n}$ as $\lambda\searrow 0$. By the usual Tauberian Theorems (see for instance [bingham1974asymptotic, Theorem A]), we obtain the result. ∎
@@ 6.1 Description of the step distribution
Recall the definitions of $\tau_{\mathsf{F}},P_{\mathsf{F}},(\tau_{\mathsf{F}}^{i},P_{\mathsf{F}}^{i})_{i\geq 0}$ at the end of Section 2.2.3. In the subcritical case $q<4$, the expected reduced length $|\overline{P_{\mathsf{F}}}|$ of $P_{\mathsf{F}}$ is infinite (this is essentially a consequence of [SheffieldScott2016QGAI, Lemma 3.7], see the second paragraph of [SheffieldScott2016QGAI, Section 3.6]). The number of $\mathsf{F}$ symbols left in the reduced word $\overline{P}_{\mathsf{F}}$ are then negligible compared to the Brownian $\sqrt{n}$ scaling. As we expect a different scaling in the critical case $q=4$, understanding the reduced length of $P_{\mathsf{F}}$ is central.
@@ 6.3 On the contribution of unmatched $\mathsf{F}$ symbols
Note that $w_{n}\log^{2}(w_{n})\sim\sqrt{n}$ as $n\to\infty$. Thus, the right-hand side goes to 0 as $B\to\infty$ uniformly over $n$ by standard simple random walk estimates (combining, say, Lévy’s inequality [VaartA.W.vander2023Wcae, Proposition A.1.2] and Hoeffding’s inequality). ∎
@@ 7.1 Tightness of the discrepancy trajectory
The main issue is that we only know how to control the discrepancy at specific random times (i.e. in the natural time parameterisation of our two explorations into the past and future). These random times are so that both explorations display a nice random walk structure, which allowed us to deduce the scaling limits. We will now write (or bound) $\mathcal{D}_{n}$ at fixed time $n$ in terms of these nicer objects. This decomposition is in some sense based on excursion theory.
@@
The process $(h^{\rightarrow}_{k}-c^{\rightarrow}_{k},k\geq 1)$ is a symmetric random walk. By Lévy’s inequality [VaartA.W.vander2023Wcae, Proposition A.1.2], we deduce that
@@
The second term is again treated using 5.11 (taking $B$ large enough), so we now focus on the first one, for fixed $B$. The process $({h^{\leftarrow}_{k}-c^{\leftarrow}_{k}},k\geq 0)$ is a symmetric random walk, so that we can use again Lévy’s inequality [VaartA.W.vander2023Wcae, Proposition A.1.2]. It gives that
@@ 7.2.1 Tightness in the Skorohod space
By [kallenberg, Theorem 23.11], to prove tightness in the Skorokhod $J_{1}$ topology, it suffices to prove tightness at a fixed time (which is given by 7.1) and that for any sequence $(\tau_{n})$ of predictable $Z^{(n)}$ stopping times, and any sequence $\varepsilon_{n}\to 0$, we have that
@@ 7.2.2 Identification of the limit
Now, the only Lévy process which has Brownian scaling is Brownian motion, as can be seen from the Lévy-Khintchine formula for stable processes [DurrettRick2019PTaE, Equation 3.7.11]. We conclude that $\beta$ has the law of (possibly correlated) planar Brownian motion $((B^{1}_{\alpha t},B^{2}_{\rho t}),t\geq 0)$ with some variance. We also know that $\alpha=1$ since $\mathcal{S}$ is a simple symmetric random walk. Finally, for any $s,t\geq 0$ $\mathbb{E}[B_{t}^{1}B_{\rho s}^{2}]=-\mathbb{E}[B_{t}^{1}B_{\rho s}^{2}]$ by symmetry of hamburgers and cheeseburgers (i.e. if the convergence holds along a subsequence for $(\mathcal{S},\mathcal{D})$, then it also holds along the subsequence for $(\mathcal{S},-\mathcal{D})$ and the limits coincide). Thus $B^{1}$ and $B^{2}$ have covariance zero, so are independent. This proves 7.3. ∎
@@
Now, the right-hand side has the law of a centred Cauchy distribution with parameter $\pi^{3}$ and the left-hand side has the law of $\sqrt{\rho}$ times a centred Cauchy distribution with parameter $\pi^{2}/2$ (either by direct calculations or using the fact that Brownian motion subordinated at an independent Brownian hitting time is Cauchy distributed, see e.g. [revuz2013continuous, Proposition III.3.11]). This implies that $\rho=4\pi^{2}$. ∎
@@
We first recall some notation and arguments from [SheffieldScott2016QGAI, Section 3.1]. Let J-{J} be the index of the first symbol in $(-\infty,0)$ whose match is in $0,+\infty)$. Notice that X(J)X(-J) is either $\mathsf{h}$ or $\mathsf{c}$, and that when $X(0)=\mathsf{F}$, we have $J=\varphi(0)$. Moreover, the reduced word $\overline{X(-J,-1)}$ can only consist of one burger production and a string of orders with the opposite type (and so importantly, no $\mathsf{F}$ symbols). Therefore, $\mathcal{S}(-J,-1)+2=|\mathcal{D}(-J,-1)|$. Furthermore, on the event that $X(0)=\mathsf{F}$, we have $\mathcal{S}(-J,-1)+1=\xi$ where $\xi$ is the step distribution of the reduced walk (Section 5.1). By (5.9), we deduce that

### v2

@@ 3.1 Exact expression of the spectral density
We first recall a few properties of the resolvent function $W$ introduced in (2.22), which are part of the so-called one-cut lemma proved in [BBG12b, Section 6] using combinatorial arguments. The basic main object we use to derive the solution of (2.23) is the spectral density $\rho$ introduced in [BBG12a], defined as follows:
@@
It is known that $\rho$ is a positive and continuous function defined on the cut $[a,b]$ that vanishes at the endpoints $a$ and $b$, and it can be deduced from [BBG12b, Equation 6.12] that $\rho$ is differentiable on (a,b)(a,b). [^7] In particular, these properties will justify the applications of the Sokhotsi–Plemelj formula below. Alternatively, we may simply assume that the Sokhotsi–Plemelj formula can be applied to $\rho$, and then after deriving our expression for $\rho$, justify this fact a posteriori by uniqueness. Recall that (2.22) implies $W(z)\sim\frac{1}{z}$ as $z\to\infty$. The Cauchy integral formula then implies that
@@
By the Sokhotski-Plemelj formula [MC11, Theorem 3.2.3], we have
@@
Then $\Phi$ is a holomorphic function in ${\mathbb{C}^{*}}\setminus[-1,1]$ whose spectral density along the cut [1,1][-1,1] is $\hat{\rho}$. By the Sokhotski-Plemelj formula [MC11, Theorem 3.2.3],
@@ 3.2 Expression and asymptotics of $F_{\ell}$
The first two claims of the above proposition are nothing but the statement of 1.4. Noting from 3.3 that the right endpoint of the cut is $b=2\sqrt{2}$, the asymptotics (3.21) match the form given by e.g. [BBG12a, Equation (5.1)], with the additional logarithmic correction that is expected from the critical case ($n=2$).
@@ 4.1 Law of a typical cluster
By Sheffield’s bijection, we may encode $\mathfrak{t}$ as a word $w$ with $\overline{w}=\emptyset$. More precisely, Sheffield’s bijection is between the word $w$ and the map $\mathfrak{t}^{\prime}$ (in one-to-one correspondence with $\mathfrak{t}$) which is obtained from $\mathfrak{t}$ by joining all the vertices on the external face of $\mathfrak{t}$ to a single extra vertex. See Section 2.4. On the other hand, conditioned on $X(0)=\mathsf{F}$ and $X(\varphi(0))=\mathsf{h}$, the envelope $\mathfrak{e}(0)$ at 0 is encoded by the $\mathsf{F}$-excursion $E:=X(\varphi(0))\cdots X(0)$. By Lemma 2.3, the probability we are after is
@@ 4.3 Exponents for the typical loop and cluster
Our next result deals with the length $|\mathfrak{L}(0)|$ of the typical loop $\mathfrak{L}(0)$, defined as the number of triangles it crosses. This is the analogue, for the critical FK model, of the first part of [BLR17, Theorem 1.1]. Note that the latter also contains a statement about the area inside $\mathfrak{L}(0)$. We should also be able to obtain such a statement from an asymptotic that we will derive in the next section, namely (5.22). However this would require some more work, and since we will not need it in the proof of our main result, we will not pursue this direction.
@@ 5.1 The step distribution of the reduced walk
We follow the strategy of [DON82, Proof of Theorem 1]. First, we relate the generating function $G(z)=\mathbb{E}[z^{\xi}]$ of $\xi$ and $r$ of (5.1). Decomposing according to the first jump, we have for all $s\in(0,1)$,
@@
One could deduce (5.9) from the fact that $\chi=2$ in Sheffield’s notation, see [SHE16, Section 3.1]. Note, however, that this would require the difficult implication of [SHE16, Lemma 3.1].
@@
We remark that 5.2 entails $F_{\xi}(\lambda)=1+o(\lambda)$ as $\lambda\to 0$. Since $\xi\geq-1$, we may use [BD74, Section 2] to deduce that $\xi$ has finite first moment and that $\mathbb{E}[\xi]=0$, which proves (5.9).
We now turn to (5.10). Using [BGT89, Corollary 8.1.7] and the Laplace transform asymptotics in 5.2, we get for free that
@@
with $a_{1}=2/\pi^{2}$. The asymptotics of the distribution function in (5.10) are more delicate and require some de Haan theory, because our case is critical (which amounts to $\alpha=1$ in the notation of [BGT89, Corollary 8.1.7]). In this case one needs to check an extra condition, which is recalled in the discussion following [BGT89, Corollary 8.1.7]. Namely, one needs to check that, if we set $\Phi(y):=y(1-{F}_{\xi}(1/y))$, then for all $r\geq 1$,
@@ 5.2 Scaling limit of the reduced walks
Recall that $\xi$ is the step distribution of $\mathtt{h}$ and $\mathtt{c}$. The asymptotics 5.2 for the Laplace transform of $\xi$ allow us to obtain the following scaling limit result for $\mathtt{h}$ (by symmetry, the same statement holds for $\mathtt{c}$).
@@
This Laplace transform is that of a $1$-stable random variable, see e.g. [SAT99, Chapter 3, Theorem 14.11]. Tracing the constants (see Lemma 5.1), this concludes the proof of 5.5. ∎
@@
We now argue that 5.6 can be seen as a scaling limit result for the primal and dual boundary length processes, when exploring towards a marked point in the infinite FK(4) random planar map. The lazy reduced walks $h^{\leftarrow}$ and $c^{\leftarrow}$ are indeed counting respectively the primal and dual boundary lengths of the component containing the triangle at 0 at each time in Sheffield’s exploration of the map, seen from infinity (see Figure 8). Whenever the lazy walk encounters a symbol $\mathsf{F}$, it jumps by an amount equal to the perimeter of the corresponding envelope, which gets disconnected from the triangle at 0. In the critical mating-of-trees coupling [AHP+23], the limiting pair $(\zeta,\zeta^{\prime})$ in 5.6 comes from subordinating planar Brownian motion on the inverse local time of excursions above the running infimum of its imaginary part. In this coupling, the two processes describe the boundary lengths in the uniform $\mathrm{CLE}_{4}$-exploration of a critical $\gamma=2$ quantum cone towards a Liouville-typical point. [^8] Technically the results of [AHP+23] are stated for quantum discs, but they are expected to extend to other types of quantum surfaces. Therefore we may interpret 5.6 as a scaling limit result for the primal and dual boundary length process in the critical FK map, to their (conjectural) continuum analogues. A convenient way to see this is to embed the portion of the map from -$\infty$ to 0 in the half-plane, as in Figure 8.
@@ 5.3 Joint law with the duration
Moreover, for $\lambda\in[\lambda_{n+1},\lambda_{n})$ we have $\psi(\lambda)\in[\psi(\lambda_{n+1}),\psi(\lambda_{n})]$ because $\eta$ is positive, so that the asymptotic (5.23) holds with $\lambda$ in place of $\lambda_{n}$ as $\lambda\searrow 0$. By the usual Tauberian Theorems (see for instance [BD74, Theorem A]), we obtain the result. ∎
@@ 6.1 Description of the step distribution
Recall the definitions of $\tau_{\mathsf{F}},P_{\mathsf{F}},(\tau_{\mathsf{F}}^{i},P_{\mathsf{F}}^{i})_{i\geq 0}$ at the end of Section 2.2.3. In the subcritical case $q<4$, the expected reduced length $|\overline{P_{\mathsf{F}}}|$ of $P_{\mathsf{F}}$ is infinite (this is essentially a consequence of [SHE16, Lemma 3.7], see the second paragraph of [SHE16, Section 3.6]). The number of $\mathsf{F}$ symbols left in the reduced word $\overline{P}_{\mathsf{F}}$ are then negligible compared to the Brownian $\sqrt{n}$ scaling. As we expect a different scaling in the critical case $q=4$, understanding the reduced length of $P_{\mathsf{F}}$ is central.
@@ 6.2 Scaling limit of the forward burger counts
$\mathbb{E}[\mathrm{e}^{-\lambda\mathcal{H}^{*}(P_{\mathsf{F}})}].=\frac{2}{\pi^{3}}\mathrm{e}^{\lambda}\int^{1}_{0}\int^{1}_{0}\log\bigg(\frac{1+\sqrt{1-u^{2}}}{u}\bigg)\log\bigg(\frac{1+\sqrt{1-v^{2}}}{v}\bigg)\frac{u(1-\pi v/2)}{u+\varepsilon}\frac{\mathrm{d}u\mathrm{d}v}{u+v}.$
@@ 6.3 On the contribution of unmatched $\mathsf{F}$ symbols
Note that $w_{n}\log^{2}(w_{n})\sim\sqrt{n}$ as $n\to\infty$. Thus, the right-hand side goes to 0 as $B\to\infty$ uniformly over $n$ by standard simple random walk estimates (combining, say, Lévy’s inequality [VW23, Proposition A.1.2] and Hoeffding’s inequality). ∎
@@ 7.1 Tightness of the discrepancy trajectory
The main issue is that we only know how to control the discrepancy at specific random times (i.e. in the natural time parametrisation of our two explorations into the past and future). These random times are so that both explorations display a nice random walk structure, which allowed us to deduce the scaling limits. We will now write (or bound) $\mathcal{D}_{n}$ at fixed time $n$ in terms of these nicer objects. This decomposition is in some sense based on excursion theory.
@@
The process $(h^{\rightarrow}_{k}-c^{\rightarrow}_{k},k\geq 1)$ is a symmetric random walk. By Lévy’s inequality [VW23, Proposition A.1.2], we deduce that
@@
The second term is again treated using 5.11 (taking $B$ large enough), so we now focus on the first one, for fixed $B$. The process $({h^{\leftarrow}_{k}-c^{\leftarrow}_{k}},k\geq 0)$ is a symmetric random walk, so that we can use again Lévy’s inequality [VW23, Proposition A.1.2]. It gives that
@@ 7.2.1 Tightness in the Skorohod space
By [KAL21, Theorem 23.11], to prove tightness in the Skorokhod $J_{1}$ topology, it suffices to prove tightness at a fixed time (which is given by 7.1) and that for any sequence $(\tau_{n})$ of predictable $Z^{(n)}$ stopping times, and any sequence $\varepsilon_{n}\to 0$, we have that
@@ 7.2.2 Identification of the limit
Now, the only Lévy process which has Brownian scaling is Brownian motion, as can be seen from the Lévy-Khintchine formula for stable processes [DUR19, Equation 3.7.11]. We conclude that $\beta$ has the law of (possibly correlated) planar Brownian motion $((B^{1}_{\alpha t},B^{2}_{\rho t}),t\geq 0)$ with some variance. We also know that $\alpha=1$ since $\mathcal{S}$ is a simple symmetric random walk. Finally, for any $s,t\geq 0$ $\mathbb{E}[B_{t}^{1}B_{\rho s}^{2}]=-\mathbb{E}[B_{t}^{1}B_{\rho s}^{2}]$ by symmetry of hamburgers and cheeseburgers (i.e. if the convergence holds along a subsequence for $(\mathcal{S},\mathcal{D})$, then it also holds along the subsequence for $(\mathcal{S},-\mathcal{D})$ and the limits coincide). Thus $B^{1}$ and $B^{2}$ have covariance zero, so are independent. This proves 7.3. ∎
@@
Now, the right-hand side has the law of a centred Cauchy distribution with parameter $\pi^{3}$ and the left-hand side has the law of $\sqrt{\rho}$ times a centred Cauchy distribution with parameter $\pi^{2}/2$ (either by direct calculations or using the fact that Brownian motion subordinated at an independent Brownian hitting time is Cauchy distributed, see e.g. [RY13, Proposition III.3.11]). This implies that $\rho=4\pi^{2}$. ∎
@@
We first recall some notation and arguments from [SHE16, Section 3.1]. Let J-{J} be the index of the first symbol in $(-\infty,0)$ whose match is in $0,+\infty)$. Notice that X(J)X(-J) is either $\mathsf{h}$ or $\mathsf{c}$, and that when $X(0)=\mathsf{F}$, we have $J=\varphi(0)$. Moreover, the reduced word $\overline{X(-J,-1)}$ can only consist of one burger production and a string of orders with the opposite type (and so importantly, no $\mathsf{F}$ symbols). Therefore, $\mathcal{S}(-J,-1)+2=|\mathcal{D}(-J,-1)|$. Furthermore, on the event that $X(0)=\mathsf{F}$, we have $\mathcal{S}(-J,-1)+1=\xi$ where $\xi$ is the step distribution of the reduced walk (Section 5.1). By (5.9), we deduce that